if the diameter of the cylindrical object is 8feet 3 inches in diameter and it it 563 feet 2 inches long how many gallons are contained in this area and what is the weight of the liquid
Answer by mananth(16949)

the diameter of the cylindrical object is 8feet 3 inches
563 feet 2 inches long


volume of cylinder = pi*r^2*h
diameter = 8ft 3 in
radius = 4ft 1.5 inch
=4 1/8 ft
=33/8 ft
...
Length = 563 feet 2 inches
=563 +1/6 ft
=3379/6 ft
Volume = pi* (33/8)^2*3379/6
= 30,104.73 cu.ft
1cu.ft = 7.48 gallons (liquid)
30,104.73 cu.ft=225,183 gallons
how many gallons are contained in this area and what is